# Computer Vision-based Adaptive Control for Back Exoskeleton Performance Optimization

### v1

@@ Abstract
Back exoskeletons can reduce musculoskeletal strain, but their effectiveness depends on support modulation and adaptive control. This study addresses two challenges: defining optimal support strategies and developing adaptive control based on payload estimation. We introduce an optimization space based on muscle activity reduction, perceived discomfort, and user preference, constructing functions to identify optimal strategies. Experiments with 12 subjects revealed optimal operating regions, highlighting the need for dynamic modulation. Based on these insights, we developed a vision-based adaptive control pipeline that estimates payloads in real-time by enhancing exoskeleton contextual understanding, minimising latency and enabling support adaptation within the defined optimisation space. Validation with 12 more subjects showed over $80\%$ accuracy and improvements across all metrics. Compared to static control, adaptive modulation reduced peak back muscle activation by up to $23\%$ while preserving user preference and minimising discomfort. These findings validate the proposed framework and highlight the potential of intelligent, context-aware control in industrial exoskeletons.
Index Terms:
Back exoskeleton, computer vision, exoskeleton optimization, deep learning, context awareness
@@ I Introduction
Musculoskeletal disorders are a growing concern in industrial workplaces, where workers in construction sites, production, logistics, and others who regularly lift heavy loads face significant risks. To address these issues, back support exoskeletons (BEs) have been developed, aiming to reduce muscular activation and spinal loads—key contributors to back impairment [1, 2, 3, 4]. Over the past few decades, advancements in the design of active BEs have improved their kinematic compatibility, reduced their weight, enhanced their acceptability, and proposed innovative designs [5, 6, 7, 8, 9]. Despite these improvements, challenges remain, particularly in terms of optimization strategies and adaptation, as current BEs lack effective mechanisms to integrate user and environmental data for optimal high-level control, leaving opportunities for significant performance gains untapped [10]. For instance, even though keeping high assistance regardless of the payload could seem a valuable strategy, since increasing the exoskeleton support level can reduce muscular effort, past studies indicate that these benefits plateau at higher support levels [11]. Furthermore, increasing support can lead to joint misalignment, discomfort, and reduced user acceptability, highlighting the need to balance multiple factors/metrics such as muscular activity, perceived discomfort, and user preference in optimizing exoskeleton performance [12, 13]. Nevertheless, introducing assistance dependencies to the payload introduces even further uncertainties on which is the best support to select. Thus, an important open question in the research deals with selecting optimal support levels based on environmental conditions or user intent to maximize performance. Another key question, in contrast, concerns the exoskeleton’s perception and adaptation capabilities. Notably, adaptive control in response to user intent and environmental factors is feasible only in active exoskeletons, which are increasingly demonstrating their adaptability and effectiveness. In particular, real-time payload estimation and adaptive support modulation have emerged as critical areas of research for enhancing exoskeleton performance [10, 14]. Furthermore, previous research highlights that the highest muscular activation in back muscles occurs during the initial phases of a lifting cycle [15], emphasizing the need for rapid and precise support modulation before the lifting process begins. Existing payload estimation techniques often fail to address this latency issue, limiting their effectiveness. Previous researchers proposed some valuable strategies to exploit exoskeleton adaptation to user needs in the payload estimation and environment comprehension enhancement attempts. One of the first and most straightforward solutions was proposed by Luis A. Mateos [16], which employed force sensing resistor insoles to measure total weight changes due to payload handling, estimate the payload, and distinguish between lifting and lowering phases. Related research by Matijevich et al. investigated a sensor network with a minimal number of sensors to accurately estimate low back loads, concluding that pressure insoles are essential for proper back load estimation as they carry crucial information about the payload [17]. Lazzaroni et al. [18, 19] and Toxiri et al. [20] employed optimized electromyography (EMG)-based armbands to measure muscular contraction while grasping objects and to modulate the exoskeleton support based on grasping strength. While this strategy demonstrated significant potential, subject-specific behaviours, persistent EMG-related drawbacks, and ergonomic considerations (e.g., skin placement requirements) limit its broader applicability. Similarly, Islam et al.[21] and Tahir et al. [22] proposed force-myography (FMG)-based sensors that correlate muscular volume changes with muscular contraction and payload estimation. Although these methods showed promising results in small-scale tests, they face challenges with output sensitivity to sensor placement and inter- and intra-subject variability, leading to low generalization capabilities [23]. Pesenti et al. [24] explored Long Short Term Memory network-based models to classify payload weights using body kinematics obtained from Inertial Measurement Units (IMU) sensors, while other researchers integrated Natural Language Processing (NLP) interfaces into exoskeleton controls to allow user-driven support modulation [25]. However, these approaches suffer from latency issues, with long inference times causing delays in response. As previously noted, the highest muscular activation occurs at the beginning of the lifting cycle, underscoring the need for exoskeletons to provide early-phase support to maximize effectiveness. To address these gaps, this manuscript explores a novel optimization framework for active back exoskeletons, considering payload estimation, muscular activity reduction, user discomfort, and preferences. We present a computer vision-based pipeline that leverages state-of-the-art techniques for object detection, selection, and payload estimation, enabling exoskeletons to operate in high-performance regions of the defined optimization space. Specifically, the contributions of this research include:
@@
  • Optimization space definition. We formulate a novel optimization space defined by two control variables, payload and assistance level, and conduct a baseline experimental campaign with 12 subjects. Using the collected data, we develop a smooth Performance Optimization Representation Function based on EMG activity reduction, user discomfort, and preference. This framework identifies areas in the optimization space that maximize exoskeleton performance.
@@
  • Vision-based payload estimation pipeline. Building on the insights from the optimization space analysis, we design a real-time computer vision pipeline to enhance environmental comprehension and payload estimation. The pipeline integrates cutting-edge deep learning models to detect objects, predict payloads, and adapt support levels according to the optimal trends found in baseline experiments.
@@
  • Integration and validation. We integrate the vision-based pipeline into the exoskeleton control loop and validate its performance through tests involving 12 subjects. The validation includes assessing EMG signals, perceived discomfort, and user preference.
@@
  • Open-source contribution. To advance the field, we provide open-source data and code as supplementary material for further scientific exploration and development.
@@
This study tackles two key challenges: defining optimal exoskeleton support regions and enabling rapid, subject-independent payload estimation. By addressing these issues, we aim to establish new performance benchmarks for active back exoskeletons in industrial settings while introducing a novel approach to enhance their contextual awareness.
@@
Outline. Section II describes the formulation of the optimization space, the development of the environmental comprehension pipeline, and the hardware used for the experiments. Section III details both baseline and validation experiments with human subject. The results are presented in Section IV and the discussion in Section V. Section VI draws the conclusions of the paper, while the Supplementary Material provides technical details on the shared dataset and code.
@@ II-A Exoskeleton Optimization Searching Space
Figure 1: Outline of the Computer Vision-Based Adaptive Control Pipeline and the architectures of the payload estimation models, where MLP stands for Multilayer Perceptron.
As introduced, in the first part of our research we aim to identify optimal operating regions for back exoskeletons by considering two variables: the assistance level ($A_{s}$) and the payload handled by the user ($P_{yl}$). The former is a controllable parameter, while the latter is determined by the environment, representing the weight of the object chosen by the user. Fig. 3-I illustrates some performance representation functions ($J_{j}$) defined within this optimization space, highlighting the exoskeleton’s optimal/suboptimal operating regions. To formulate the optimization functions, we conducted experiments with 12 subjects, sampling discrete points in the space as described in section III-A, resulting in a dataset:
@@
where $x^{i}=(A_{s}^{i},P_{yl}^{i})$ are pairs of assistance and payload levels, and $t^{i}$ represent samples of the objective function ($J_{j}^{i}$). Using Gaussian Processes (GP) with a Radial Basis Function (RBF) kernel implemented via the GPy library in Python [26], we interpolate $D$ to obtain a smooth, continuous function [27]:
@@
The Total Performance Optimization Representation Function (T-ORF) combines three contributions: i) EMG activity reduction optimization representation function (EMG-ORF) $J_{EMG}(A_{s},P_{yl})$, ii) perceived discomfort optimization representation function (D-ORF) $J_{dsc}(A_{s},P_{yl})$, and iii) user preference optimization representation function (P-ORF) $J_{prf}(A_{s},P_{yl})$. During each test (see Section III-A), $J_{EMG}$ measures the percentage reduction in muscular activity when using the exoskeleton compared to no-exoskeleton condition. For $J_{dsc}$, a questionnaire with the following nine weighted questions ($q_{i},:i=1,...9$) assesses discomfort:
@@
    I feel highly comfortable in the current setting.
@@
    My fatigue is reduced in the current setting.
@@
    I am satisfied with the current setting.
@@
    The exoskeleton is useful for my task in the current setting.
Questions are rated on a scale from 1 (strongly disagree) to 5 (strongly agree). The weighting scheme prioritizes questions that are more directly related to discomfort (i.e., Questions 1, 2, and 4) by assigning them higher weights, while questions with a more indirect connection to discomfort receive lower weights. Specifically, the weights used in the formulation of the discomfort function were as follows: $w_{1}=2$, $w_{2}=2$, $w_{3}=1.5$, $w_{4}=2$, $w_{5}=1$, $w_{6}=1$, $w_{7}=1$, $w_{8}=1$, and $w_{9}=1$. This reflects a deliberate emphasis on the most relevant factors influencing perceived discomfort. With scores flipped for positively framed questions (2, 7, 8, 9). The final sample of the discomfort function is calculated as:
$J_{dsc}(A_{s}^{i},P_{yl}^{i})=\sum_{k=1}^{9}w_{i}q_{i}.$ (3)
For $J_{prf}$, subjects indicate their preferred assistance between light and strong when lifting light ($5$ kg) and heavy ($15$ kg) weights. After GP interpolation and normalization, the total function is:
@@
The $J_{EMG}$ component is weighted higher as it provides a quantitative performance measure, while $J_{dsc}$ and $J_{prf}$ are qualitative. The total function $J_{tot}(A_{s},P_{yl})$ reflects the negative cost of operating in a specific region, with higher values indicating better performance. This optimization space guides the selection of assistance levels to maximize overall performance based on muscular activity reduction, perceived discomfort, and user preference.
@@ II-B High-Level Control Strategy for BE Optimization
After formulating the optimization space as outlined in section II-A, our results indicate that a static support level cannot maximize performance across all payload conditions. Instead, the back exoskeleton must adapt its control strategy to the payload. To address this, we developed a Computer Vision-based Adaptive Control Pipeline (CVAC) for high-level support adaptation, illustrated in Fig. 1 (a). This pipeline integrates RGB images from an embedded camera for object detection and depth data from a depth sensor for object selection, identifying the most likely object to be picked up by the user. The selected object’s image is processed by a fine-tuned Vision Transformer (ViT) [28] to estimate object weight directly, classifying it as light, medium, or heavy. This classification guides the selection of low, medium, or high assistance levels based on the optimization space defined in section II-A. This work focuses on the general case where the weight of objects (e.g., transparent boxes) is unknown, requiring a deep learning model to infer weight based on object content. The following sections describe each layer of the high-level control pipeline in detail.
@@ II-B1 Object Detection & Selection
The first stage of the pipeline handles object detection and selection. To enhance environmental awareness, we fine-tuned YOLOv11 [29] on a custom dataset to detect transparent boxes, which allows subsequent layers to infer weight from the box’s contents. The dataset comprises 831 images acquired at the BioRobotics Lab at the National University of Singapore (NUS), and labeled with “box” class and bounding box coordinates, split into 88% training, 9% validation, and 3% testing. The fine-tuned YOLO operates on cloud GPUs in real-time, processing RGB frames from the embedded camera. Since YOLO may detect multiple objects, a Time of Flight (ToF) distance map sensor estimates each object’s approximate distance information, which we use to identify the one most likely to be picked up. Although we tested state-of-the-art Monocular Depth Estimation (MDE) models such as DepthAnythingv2 [30] and MiDaS [31] fine-tuned for metric depth estimation (further details in supplementary materials), we opted for the Arducam ToF camera due to its superior performance in indoor environments within our application range ([04]m[0-4]m). Using the distance map data, we compute a probability factor $\alpha_{i}$ for each detected object as follows:
@@
where $\lambda_{\theta}=0.1$ and $\lambda_{d}=0.5$ control decay rates, optimized during preliminary tests, $\theta_{i}$ is the angle of the $i$-th object in the camera’s field of view (with 0 at the center), $d_{i}$ is the object’s distance, and 10 is just a scale factor. Objects farther from the user or closer to the field’s boundaries are assigned lower probabilities. After normalizing the factors using softmax, we select the object with the highest pick-up probability as a candidate. If the object with highest probability is within $2m$, we consider the lifting cycle starting soon, we block the recognition pipeline, we crop the environment image on the bounding box of the candidate object, and pass it to the next layer.
@@ II-B2 Payload Estimation
After the object detection and selection layer (section II-B1), we feed the selected object’s cropped image into a fine-tuned deep-learning model for direct weight classification between light, medium, and heavy. For this purpose, we compare two state-of-the-art models: MobileNetv2 [32], a Convolutional Neural Network (CNN) optimized for real-time performance, and DINOv2, a ViT-based foundation model well known for high-quality feature extraction and contextual understanding capabilities [33, 34, 35].
MobileNetv2 Approach (Fig. 1, b): MobileNetv2 extracts features from input images of shape $(256,256,3)$, outputting $(1,8,8,1024)$ feature maps, which are flattened into 65,536 features. These are mapped to a 32-dimensional space using a four-layer fully connected network. To enhance physical perception, we concatenate physical features (bounding box dimensions and object distance) with the extracted vector, resulting in a 35-dimensional input to a final linear layer for weight classification. We use ReLU activations, softmax for output, RMSprop optimizer ($lr=1e^{-4}$), and categorical cross-entropy loss. In this case, we only use MobileNetv2 to extract features from images, and we train the layers defined above on the new latent space of extracted features.
DINOv2 Approach (Fig. 1, c): We use DINOv2 pretrained base model available at [36] to process the input image, and we stuck a single-head attention block at the end of it. The output features $(1,324,768)$ are averaged, producing a 768-dimensional vector fed into an MLP with dropout (0.4) and ReLU activation, and softmax for output. Unlike MobileNetv2, because DINOv2 showed strong depth perception capabilities in past research, no physical features are added in this case. We use Adam optimizer ($lr=1e^{-4}$) and categorical cross-entropy loss. In this case, we fine-tune DINOv2 to our specific classification objective, freezing its core weights while updating the parameters of the last additional layers during training.
In both cases, the models were trained on a dataset of 5480 labeled images (32.03%, 34.01%, and 33.96% distributions for light, medium, and heavy weights) collected at the BioRobotics Lab at the National University of Singapore (NUS). The first model was trained on features from MobileNetv2 for 40 epochs with a batch size of 400, while DINOv2 was fine-tuned for 20 epochs with a batch size of 40. Codes, datasets, and further details on the choice of hyperparameters are provided as supplementary materials.

### v2

# Toward Context-Aware Exoskeleton Assistance: Integrating Computer Vision Payload Estimation with a User-Centric Optimization Space

@@ Abstract
Back-support exoskeletons (BSEs) mitigate musculoskeletal strain, yet their efficacy depends on precise, context-aware modulation. This paper introduces a user-centric optimization framework and a vision-based adaptive control strategy for industrial BSEs. First, we constructed a multi-metric optimization space, integrating electromyography reduction, perceived discomfort, and user preference, through baseline experiments with 12 subjects. This revealed a non-linear relationship between optimal assistance and payload. Second, we developed a predictive computer vision pipeline using a Vision Transformer (DINOv2) to estimate payloads before lifting, effectively overcoming actuation latency. Validation with 12 subjects confirmed the system’s robustness, achieving over 82% estimation accuracy. Crucially, the adaptive controller reduced peak back muscle activation by up to 23% compared to static baselines while optimizing user comfort. These results validate the proposed framework, demonstrating that pre-lift environmental perception and user-centric optimization significantly enhance physical assistance and human-robot interaction in industrial settings.
@@ I Introduction
Musculoskeletal disorders are a growing concern in industrial workplaces, particularly in construction, production, and logistics, where workers regularly lift heavy loads. To address these issues, back-support exoskeletons (BSEs) have been developed to reduce muscular activation and spinal loads, which are key contributors to back impairment [40, 26, 2, 9]. Over the past few decades, design advancements have improved active BSEs in terms of kinematic compatibility, weight, and acceptability [11, 5, 34, 37, 17]. Despite these improvements, challenges remain regarding optimization and adaptation; current BSEs lack effective mechanisms to integrate user and environmental data for high-level control, leaving significant performance gains untapped [7]. For instance, although maintaining high assistance regardless of payload may seem valuable for reducing muscular effort, studies indicate that benefits plateau at higher support levels [43]. Furthermore, increasing support can cause joint misalignment, discomfort, and reduced acceptability, highlighting the need to balance muscular activity, perceived discomfort, and user preference [21, 15]. Introducing assistance dependencies on payload adds further uncertainty regarding the best support selection. Thus, selecting optimal support levels based on environmental conditions or user intent remains an open research question. A contrasting but key challenge concerns the exoskeleton’s perception. Adaptive control is feasible only in active exoskeletons, where real-time payload estimation and adaptive modulation are critical for enhancing performance [7, 1]. Notably, the highest back muscle activation occurs during the initial lifting phase [41], emphasizing the need for rapid support modulation, but existing payload estimation techniques often fail to address this latency, limiting their effectiveness.
Previous researchers have proposed strategies to exploit adaptation for payload estimation and environmental comprehension. One of the first solutions by Mateos [22] employed force sensing resistor insoles to measure weight changes and distinguish lifting phases. Related research by Matijevich et al. [23] concluded that pressure insoles are essential for proper back load estimation. Lazzaroni et al. [19], [20] and Toxiri et al. [38] used optimized electromyography (EMG) armbands to modulate support based on grasping strength. While potentially significant, this strategy is limited by subject-specific behaviors, EMG drawbacks, and ergonomic considerations. Similarly, Islam et al. [39] and Tahir et al. [36] proposed force-myography (FMG) sensors, though these face challenges with sensor placement sensitivity and generalization [44]. Pesenti et al. [28] explored LSTM models using IMU kinematics, while others integrated Natural Language Processing (NLP) for user-driven modulation [24]. However, these approaches often suffer from latency and long inference times. As noted, since peak activation occurs at the lifting cycle’s start, early-phase support is crucial.
To address these gaps, this manuscript explores a novel optimization framework for active back-support exoskeletons, considering payload estimation, muscular activity reduction, user discomfort, and preferences. We then present a computer vision-based pipeline leveraging state-of-the-art object detection and payload estimation to enable operation in high-performance regions. Specifically, the contributions include:
@@
  • Optimization space definition. We formulate a novel optimization space defined by payload and assistance level, conducting a baseline campaign with 12 subjects. Using this data, we develop a smooth Performance Optimization Representation Function based on EMG reduction, discomfort, and preference to identify performance-maximizing regions.
@@
  • Vision-based payload estimation pipeline. Building on the optimization analysis, we design a real-time computer vision pipeline to enhance environmental comprehension. The pipeline integrates deep learning models to detect objects, predict payloads, and adapt support levels according to the optimal trends found.
@@
  • Integration and validation. We integrate the vision-based pipeline into the exoskeleton control loop and validate its performance with 12 subjects, assessing EMG signals, perceived discomfort, and user preference.
@@
  • Open-source contribution. We provide open-source data and code as Supplementary Material to facilitate further scientific exploration.
@@
This study tackles two key challenges: defining optimal support regions and enabling rapid, subject-independent payload estimation. By addressing these, we aim to establish new benchmarks for active back-support exoskeletons in industrial settings while enhancing contextual awareness.
@@
Outline. Section II describes the optimization formulation, environmental pipeline, and hardware. Section III details baseline and validation experiments. Results and discussion are presented in Section IV and Section V, followed by conclusions in Section VI.
@@ II-A Exoskeleton Optimization Searching Space
Figure 1: Overview of the CVAC pipeline (a): images captured by the RPI are transmitted to the cloud workstation for object detection and payload estimation, and the resulting predictions are returned to the RPI for serial communication with the exoskeleton. Panels (b) and (c) illustrate the architectures of the payload-estimation models; MLP denotes Multilayer Perceptron.
As introduced, in the first part of our research, we aim to identify optimal operating regions for back-support exoskeletons using two variables: assistance level ($A_{s}$) and user payload ($P_{yl}$). $A_{s}$ is a controllable parameter, while $P_{yl}$ is determined by the environment. Fig. 4-I illustrates performance representation functions ($J_{j}$) within this space, highlighting optimal operating regions. To formulate these functions, we sampled discrete points with 12 subjects (see Section III-A) to generate a dataset:
@@
where $x^{i}=(A_{s}^{i},P_{yl}^{i})$ are assistance-payload pairs, $t^{i}$ are objective function samples ($J_{j}^{i}$), and $N$ represents the number of samples in the dataset. Using Gaussian Processes (GP) with a Radial Basis Function (RBF) kernel [14, 32], we fit $D$ to obtain continuous functions:
@@
The Total Performance Optimization Representation Function (T-ORF) combines three components: i) EMG activity reduction ($J_{EMG}$), ii) perceived discomfort ($J_{dsc}$), and iii) user preference ($J_{prf}$). $J_{EMG}$ measures the percentage reduction in muscular activity compared to the no-exoskeleton condition. For $J_{dsc}$, we assess discomfort with a questionnaire with nine weighted questions ($q_{i}$):
@@
    I feel highly comfortable in the current setting. -⃝-⃝
@@
    My fatigue is reduced in the current setting. -⃝-⃝
@@
    I am satisfied with the current setting. -⃝-⃝
@@
    The exoskeleton is useful for my task in the current setting. -⃝-⃝
Questions are rated from 1 (strongly disagree) to 5 (strongly agree). Weights prioritize factors directly related to discomfort ($w_{1,2,4}=2$, $w_{3}=1.5$) over indirect connections ($w_{5-9}=1$). Scores are flipped for positively framed questions (2, 7, 8, 9, highlighted with a -⃝-⃝ sign). The final discomfort sample is:
$J_{dsc}(A_{s}^{i},P_{yl}^{i})=\sum_{k=1}^{9}w_{k}q^{i}_{k}.$ (3)
For $J_{prf}$, subjects indicate their preferred assistance (light vs. strong) when lifting light ($5$ kg) and heavy ($15$ kg) weights. After interpolation and normalization, the total function is:
@@
We weight $J_{EMG}$ highest as it provides a quantitative measure, while we weight qualitative metrics ($J_{dsc},J_{prf}$) lower. The total function $J_{tot}$ reflects the negative cost of operating in a region, with higher values indicating better performance, guiding the selection of optimal assistance levels.
@@ II-B High-Level Control Strategy for BE Optimization
Following the optimization formulation in Section II-A, our results indicate that static support fails to maximize performance across varying payloads. Consequently, the back-support exoskeleton requires payload-adaptive control. To achieve this, we developed the Computer Vision-based Adaptive Control (CVAC) pipeline (Fig. 1-a). This pipeline integrates RGB and depth data to detect and select the target object. The resulting image is processed by a fine-tuned Vision Transformer (ViT) [12] to classify the weight (light, medium, or heavy), determining the optimal assistance level. This approach specifically addresses scenarios where object weight is unknown (e.g., transparent boxes), requiring inference based on visual content. The pipeline components are detailed below.
@@ II-B1 Object Detection & Selection
The first stage of the pipeline performs object detection and selection. To enhance environmental awareness, we fine-tuned YOLOv11 [33] on a custom dataset of 831 images collected at the BioRobotics Lab at the National University of Singapore (NUS), labeled with a “box” class and bounding box coordinates, and split into 88% training, 9% validation, and 3% testing. The fine-tuned model runs on cloud GPUs in real-time, processing RGB frames from the embedded camera. Since YOLO may detect multiple objects, a Time-of-Flight (ToF) distance sensor provides approximate depth estimates, which we use to identify the object most likely to be picked up. Although we tested state-of-the-art Monocular Depth Estimation (MDE) models such as DepthAnythingv2 [46] and MiDaS [31] fine-tuned for metric depth estimation (see Supplementary Materials), the Arducam ToF camera offered superior performance indoors within our operational range ($0\!-\!4]\,\mathrm{m}$). Using the distance map, we compute a probability factor $\alpha_{i}$ for each detected object:
@@
where $\lambda_{\theta}=0.1$ and $\lambda_{d}=0.5$ control decay rates (optimized during preliminary tests), $\theta_{i}$ is the angle of the $i$-th object in the camera’s field of view (with 0 at the center), $d_{i}$ is the object’s distance, and 10 is a scale factor. We assign lower probabilities to objects farther from the user or near the field boundaries. After softmax normalization, we select the object with the highest pick-up probability. If this object is within $2\,\mathrm{m}$, we consider the lifting cycle imminent, block the recognition pipeline, crop the environment image to its bounding box, and pass it to the next layer.
@@ II-B2 Payload Estimation
After the object detection and selection layer (Section II-B1), we feed the cropped image of the selected object into a fine-tuned deep-learning model for direct weight classification (light, medium, heavy). We compare two state-of-the-art architectures: MobileNetv2 [35], a CNN optimized for real-time performance, and DINOv2, a ViT-based foundation model known for high-quality feature extraction, contextual understanding, and strong transfer learning capabilities [45, 47, 25, 6, 10].
MobileNetv2 Approach (Fig. 1, b): MobileNetv2 processes input images of size $(256,256,3)$ and outputs feature maps of shape $(1,8,8,1024)$, which we flatten into 65,536 features. We project these into a 32-dimensional latent space through a four-layer fully connected network. To integrate physical information, we concatenate bounding box dimensions and object distance with the extracted vector, yielding a 35-dimensional input to a final linear classifier. We use ReLU activations, softmax output, RMSprop optimizer ($lr=1e^{-4}$), and categorical cross-entropy loss. In this configuration, MobileNetv2 serves as a frozen feature extractor, and we train only the subsequent layers for classification on the latent features extracted.
DINOv2 Approach (Fig. 1, c): We use the pretrained DINOv2 base model [27] and append a single-head attention block. We average the output features of shape $(1,324,768)$ into a 768-dimensional vector and pass it to an MLP with dropout (0.4), ReLU activation, and a softmax classifier. Since DINOv2 has demonstrated strong depth-perception capabilities in prior work, we don’t add any additional physical features. We use the Adam optimizer for training ($lr=1e^{-4}$) with categorical cross-entropy loss. Here, we fine-tune DINOv2 by freezing its core weights and updating only the appended layers.
We trained both models on a dataset of 5480 labeled images (32.03%, 34.01%, and 33.96% for light, medium, and heavy), collected at the BioRobotics Lab at the National University of Singapore (NUS). We trained the MobileNetv2-based classifier for 40 epochs with a batch size of 400, whereas we fine-tuned DINOv2 for 20 epochs with a batch size of 40. We provide codes, datasets, and additional details on hyperparameter choices in the Supplementary Materials.